# Spherical maximal functions and Hardy spaces for Fourier integral operators

### v2

@@ 2.3. Spaces on manifolds
Up to norm equivalence, the definition of $\mathcal{H}^{s,p}_{FIO}(M)$ is independent of the choice of $K$ and $(\psi_{\kappa})_{\kappa\in K}$ as bove. This follows from a simple calculation, using (2.8) and the invariance of $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ under coordinate changes and multiplication by smooth cutoffs (see also [20, Proposition 4.6]).
@@
Since $K$ is finite, one could equally well replace the $\ell^{p}$ norms in (2.9) and (2.10) by an $\ell^{q}$ norm for any $1\leq q\leq\infty$. On the other hand, one can define $\mathcal{H}^{s,p}_{FIO}(M)$ on general Riemannian manifolds with bounded geometry, a class which includes both compact manifolds and $\mathbb{R}^{n}$, and for this larger class the $\ell^{p}$ norm in (2.9) is natural, as is already evident on $\mathbb{R}^{n}$ (see [20, Theorem 3.6]).
@@
The basic properties of $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ carry over to $\mathcal{H}^{s,p}_{FIO}(M)$, for all $p\in[1,\infty]$ and $s\in\mathbb{R}$. For example, $\mathcal{H}^{s,p}_{FIO}(M)$ is invariant under Fourier integral operators of order zero associated with a local canonical graph, cf. [20, Theorem 4.16], and the Sobolev embeddings
@@
hold, by [20, Theorem 4.15]. Here $\mathcal{H}^{s,p}(M)$ is defined in the same way as $\mathcal{H}^{s,p}_{FIO}(M)$, in local coordinates (in particular, (2.11) then follows directly from (2.2)). We also note that $\mathcal{D}(M,\Omega_{1/2})$ lies dense in $\mathcal{H}^{s,p}_{FIO}(M)$ if $p<\infty$.
@@ 3.1. Fourier integral operators
We first prove a theorem about maximal functions for the Fourier integral operators from Proposition 2.2.
@@
Now, to prove (3.1), we first deal with the low-frequency component of $a$. Let $r\in C^{\infty}_{c}(\mathbb{R}^{n})$, and note that $e^{it\phi(D)}a(tD)r(D)$ has kernel $K_{t}$ given by
$K_{t}(x,y)=\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}e^{i((x-y)\cdot\eta+t\phi(\eta))}a(t\eta)r(\eta)\mathrm{d}\eta,$
for $t\in I$ and $x,y\in\mathbb{R}^{n}$. So integration by parts yields, for each $N\geq 0$, a $C_{N}\geq 0$ such that $|x-y|^{N}|K_{t}(x,y)|\leq C_{N}$. Let $r^{\prime}\in C^{\infty}_{c}(\mathbb{R}^{n})$ be such that $r(D)=r(D)r^{\prime}(D)$. Since $r^{\prime}(D)$ is smoothing, Young’s inequality and (2.2) yield
@@
$\lesssim\int_{\mathbb{R}^{n}}\Big{(}\int_{\mathbb{R}^{n}}(1+|x-y|)^{-N}|(r^{\prime}(D)f)(y)|\mathrm{d}y\Big{)}^{p}\mathrm{d}x\lesssim\|r^{\prime}(D)f\|_{L^{p}(\mathbb{R}^{n})}^{p}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}^{p}$
for $N\geq 0$ large enough. In the remainder, it thus suffices to show that
@@
Next, note that
@@
It remains to extend (3.1) to general $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Here it is relevant to note that, a priori, it is not even clear why the left-hand side of (3.1) is well defined for such $f$. Indeed, initially $e^{it\phi(D)}a(tD)f$ is defined as a tempered distribution for each $t\in I$, but Proposition 2.2 shows that it can be identified with an element of $L^{p}(\mathbb{R}^{n})$. We have to prove that there exists a set $U\subseteq\mathbb{R}^{n}$ of measure zero such that $x\mapsto\sup_{t\in I}|e^{it\phi(D)}a(tD)f(x)|\mathbf{1}_{\mathbb{R}^{n}\setminus U}(x)$ is measurable and satisfies (3.1).
@@
(3.5) $\big{\|}\sup_{t\in I}|F(\cdot,t)|\big{\|}_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}.$
We may also suppose that there exists a set $U\subseteq\mathbb{R}^{n}$ of measure zero such that, for all $x\notin U$, one has $F_{k}(x,t)\to F(x,t)$, uniformly in $t\in I$.
On the other hand, by Proposition 2.2 (1), for each $t\in I$ one has $F_{k}(\cdot,t)\to e^{it\phi(D)}a(tD)f(\cdot)$ in $L^{p}(\mathbb{R}^{n})$ as $k\to\infty$. So there exists a set $U_{t}\subseteq\mathbb{R}^{n}$ of measure zero, and a subsequence $(F_{m_{k,t}})_{k=0}^{\infty}$ of $(F_{k})_{k=0}^{\infty}$, such that $F_{m_{k,t}}(x,t)\to e^{it\phi(D)}a(tD)f(x)$ for $x\notin U_{t}$. Hence $e^{it\phi(D)}a(tD)f(x)=F(x,t)$ for $x\notin U\cup U_{t}$. Since $e^{it\phi(D)}a(tD)f$ is only determined up to sets of measure zero, we can choose a representative such that $e^{it\phi(D)}a(tD)f(x)\mathbf{1}_{\mathbb{R}^{n}\setminus U}(x)=F(x,t)$ for all $x\in\mathbb{R}^{n}$. Now apply (3.5). ∎
@@
for any compact interval $I\subseteq(0,\infty)$. Moreover, if $f\in\overline{\mathcal{S}(\mathbb{R}^{n})}\subseteq\mathcal{H}^{s,\infty}_{FIO}(\mathbb{R}^{n})$, then $t\mapsto e^{it\phi(D)}a(tD)f$ is continuous as an $L^{\infty}(\mathbb{R}^{n})$-valued map, and (3.6) holds as well. Finally, by Theorems 4.2 and 4.5, (3.7) is essentially sharp under the same assumptions as mentioned below Theorem 3.1.
The case where $p=\infty$ in Theorem 3.4 and Propositions 3.11 and 3.14 below is straightforward for similar reasons.
@@
The proof is almost identical to that of Theorem 3.1. Firstly, (3.9) follows from (3.8), and it suffices to prove (3.8) for $f\in\mathcal{S}(\mathbb{R}^{n})$, just as in the proof of Theorem 3.1. For such $f$ the left-hand side of (3.8) is well-defined.
Now, to prove (3.8), the low-frequency component $T_{t}r(D)$ of each $T_{t}$ is a smoothing operator for $r\in C^{\infty}_{c}(\mathbb{R}^{n})$. One can therefore apply Young’s inequality and (2.2) to this term. On the other hand, the arguments from the proof of Theorem 3.1 and the bounds in Proposition 2.3 can be combined to deal with the high frequencies. ∎
@@ 3.2. Hypersurfaces
In this subsection we use Theorems 3.1 and 3.4 to prove results about maximal functions associated with hypersurfaces in $\mathbb{R}^{n}$.
For $\Sigma\subseteq\mathbb{R}^{n}$ a (smooth) hypersurface, $\psi\in C^{\infty}_{c}(\Sigma)$, $f\in C(\mathbb{R}^{n})$ and $x\in\mathbb{R}^{n}$, set
@@
Here and throughout, $\mathrm{d}\sigma$ is the induced Lebesgue measure on $\Sigma$. The following result contains Theorem 1.1 as a special case. By Theorem 4.8, it is essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty]$ whenever $\psi$ is not identically zero.
@@
Let $\Sigma\subseteq\mathbb{R}^{n}$ be a hypersurface with non-vanishing Gaussian curvature, and let $\psi\in C^{\infty}_{c}(\Sigma)$. Then, for all $p\in[1,\infty]$ and $s>d(p)+\frac{1}{p}-\frac{n-1}{2}$, there exists a $C\geq 0$ such that
@@
Fix $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Then the left-hand side of (3.13) is well defined. Moreover, $\mathcal{A}_{t}f=b(tD)f$, where $b:=\widehat{\psi\mathrm{d}\sigma}\in C^{\infty}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ and $b(tD)f$ is initially defined as a tempered distribution. In fact, after decomposing the support of $\psi$ into finitely many pieces and changing coordinates, we may suppose that
$b(\xi)=a(\xi)e^{i\phi(\xi)}+e(\xi)$
for all $\xi\in\mathbb{R}^{n}$ (see e.g. [36, Section VIII.5.7] or [33, Theorem 1.2.1]). Here $\phi\in C^{\infty}(V)$ is real-valued and positively homogeneous of degree $1$, defined on an open conic subset $V\subseteq\mathbb{R}^{n}\setminus\{0\}$, and $\mathrm{rank}\,\partial_{\xi\xi}^{2}\phi(\xi)=n-1$ for all $\xi\in V$. Also, $a\in S^{-(n-1)/2}(\mathbb{R}^{n})$ satisfies $\mathrm{supp}(a)\subseteq V^{\prime}\cup\{0\}$ for some conic $V^{\prime}\subseteq V$ which is closed in $\mathbb{R}^{n}\setminus\{0\}$, and $e\in S^{m}(\mathbb{R}^{n})$ for all $m\in\mathbb{R}$. It follows that
$\mathcal{M}_{\Sigma,\psi}f(x)\leq\sup_{1\leq t\leq 2}|e^{it\phi(D)}a(tD)f(x)|+|e(tD)f(x)|$
for all $x\in\mathbb{R}^{n}$.
Now, for each $s>d(p)+\frac{1}{p}-\frac{n-1}{2}$, Theorem 3.1 and Remark 3.3 yield
@@
On the other hand, by (3.10) and Proposition 2.2 (both with $\phi\equiv 0$ and $m=d(p)-s(p)-\frac{n-1}{2}$), one has

### v3

@@ 2.3. Spaces on manifolds
Up to norm equivalence, the definition of $\mathcal{H}^{s,p}_{FIO}(M)$ is independent of the choice of $K$ and $(\psi_{\kappa})_{\kappa\in K}$ as bove. This follows from a simple calculation, using (2.8) and the invariance of $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ under coordinate changes and multiplication by smooth cutoffs (see also [22, Proposition 4.6]).
@@
Since $K$ is finite, one could equally well replace the $\ell^{p}$ norms in (2.9) and (2.10) by an $\ell^{q}$ norm for any $1\leq q\leq\infty$. On the other hand, one can define $\mathcal{H}^{s,p}_{FIO}(M)$ on general Riemannian manifolds with bounded geometry, a class which includes both compact manifolds and $\mathbb{R}^{n}$, and for this larger class the $\ell^{p}$ norm in (2.9) is natural, as is already evident on $\mathbb{R}^{n}$ (see [22, Theorem 3.6]).
@@
The basic properties of $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ carry over to $\mathcal{H}^{s,p}_{FIO}(M)$, for all $p\in[1,\infty]$ and $s\in\mathbb{R}$. For example, $\mathcal{H}^{s,p}_{FIO}(M)$ is invariant under Fourier integral operators of order zero associated with a local canonical graph, cf. [22, Theorem 4.16], and the Sobolev embeddings
@@
hold, by [22, Theorem 4.15]. Here $\mathcal{H}^{s,p}(M)$ is defined in the same way as $\mathcal{H}^{s,p}_{FIO}(M)$, in local coordinates (in particular, (2.11) then follows directly from (2.2)). We also note that $\mathcal{D}(M,\Omega_{1/2})$ lies dense in $\mathcal{H}^{s,p}_{FIO}(M)$ if $p<\infty$.
@@ 3.1. Fourier integral operators
To deal with the low-frequency components of the operators that appear in this subsection, we will use the following quantified version of [10, Lemma 1.17].
Lemma 3.1.
There exists a $C\geq 0$ such that the following holds. Let $b\in C^{\infty}(\mathbb{R}^{n}\times(\mathbb{R}^{n}\setminus\{0\}))$ be such that $b(x,\eta)=0$ for all $(x,\eta)\in\mathbb{R}^{2n}$ with $|\eta|>1$, and such that
$C_{\alpha}:=\sup\{|\eta|^{\max(0,-1+|\alpha|)}|\partial_{\eta}^{\alpha}b(x,\eta)|\mid(x,\eta)\in\mathbb{R}^{n}\times(\mathbb{R}^{n}\setminus\{0\})\}<\infty$
for every $\alpha\in\mathbb{Z}_{+}^{n}$ with $|\alpha|\leq n+1$. Set $C_{b}:=\max\{C_{\alpha}\mid\alpha\in\mathbb{Z}_{+}^{n},|\alpha|\leq n+1\}$. Then
$\Big|\int_{\mathbb{R}^{n}}e^{i(x-y)\cdot\eta}b(x,\eta)\mathrm{d}\eta\Big|\leq CC_{b}\langle x-y\rangle^{-(n+1)}(1+\log(1+|x-y|))$
for all $x,y\in\mathbb{R}^{n}$.
Proof.
The proof is as in [10, Lemma 1.17]. For the convenience of the reader, we include the argument here.
Due to the support condition on $b$, it suffices to consider $x,y\in\mathbb{R}^{n}$ with $|x-y|\geq 1$. Set
$\beta(x,y,\eta):=(-i|x-y|^{-1}(x-y)\cdot\partial_{\eta})^{n}b(x,\eta)$
for $\eta\in\mathbb{R}^{n}\setminus\{0\}$, and let $\chi\in C^{\infty}_{c}(\mathbb{R}^{n})$ be such that $\chi(\eta)=1$ if $|\eta|\leq 1$, and $\chi(\eta)=0$ if $|\eta|\geq 2$. Integration by parts then allows us to write
$|x-y|^{n}\int_{\mathbb{R}^{n}}e^{i(x-y)\cdot\eta}b(x,\eta)\mathrm{d}\eta=\int_{\mathbb{R}^{n}}e^{i(x-y)\cdot\eta}\beta(x,y,\eta)\mathrm{d}\eta$
$=\int_{\mathbb{R}^{n}}e^{i(x-y)\cdot\eta}\chi(\eta/\varepsilon)\beta(x,y,\eta)\mathrm{d}\eta+\int_{\mathbb{R}^{n}}e^{i(x-y)\cdot\eta}(1-\chi(\eta/\varepsilon))\beta(x,y,\eta)\mathrm{d}\eta,$
for a given $0<\varepsilon<1$.
By integrating in spherical coordinates, one sees that the first term on the second line above is bounded in absolute value by a constant multiple of $C_{b}\varepsilon$. On the other hand, integrating by parts one more time shows that the second term equals
$-|x-y|^{-1}\int_{\mathbb{R}^{n}}e^{i(x-y)\cdot\eta}\varepsilon^{-1}(-i|x-y|^{-1}(x-y)\cdot\partial_{\eta}\chi)(\eta/\varepsilon)\beta(x,y,\eta)\mathrm{d}\eta$
$+|x-y|^{-1}\int_{\mathbb{R}^{n}}e^{i(x-y)\cdot\eta}(1-\chi(\eta/\varepsilon))(-i|x-y|^{-1}(x-y)\cdot\partial_{\eta}\beta(x,y,\eta))\mathrm{d}\eta.$
Then integration in spherical coordinates shows that the term on the first line is bounded in absolute value by a constant multiple of $C_{b}|x-y|^{-1}$, and that the second term is bounded in absolute value by a multiple of $C_{b}|x-y|^{-1}\log\varepsilon^{-1}$. Finally, setting $\varepsilon=|x-y|^{-1}$ concludes the proof. ∎
We are now ready to prove a theorem about maximal functions for the Fourier integral operators from Proposition 2.2.
@@
The low frequencies
Now, to prove (3.1), we first deal with the low-frequency component of $a$. Let $r\in C^{\infty}_{c}(\mathbb{R}^{n})$ be such that $r(\eta)=1$ if $|\eta|\leq 1/2$, and $r(\eta)=0$ if $|\eta|>1$. Note that $e^{it\phi(D)}a(tD)r(D)$ has kernel $K_{t}$ given by
$K_{t}(x,y)=\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}e^{i((x-y)\cdot\eta}e^{it\phi(\eta)}a(t\eta)r(\eta)\mathrm{d}\eta,$
for $t\in I$ and $x,y\in\mathbb{R}^{n}$. Moreover, $\eta\mapsto e^{it\phi(\eta)}a(t\eta)r(\eta)$ is smooth on $\mathbb{R}^{n}\setminus\{0\}$, and
$\sup\big\{|\eta|^{\max(0,-1+|\alpha|)}\partial_{\eta}^{\alpha}\big(e^{it\phi(\eta)}a(t\eta)r(\eta)\big)\,\big|\,t\in I,\eta\in\mathbb{R}^{n}\setminus\{0\}\big\}<\infty$
for all $\alpha\in\mathbb{Z}_{+}^{n}$. Hence
$\sup\{\langle x-y\rangle^{n+1/2}|K_{t}(x,y)|\mid x,y\in\mathbb{R}^{n},t\in I\}<\infty,$
by Lemma 3.1.
Next, let $r^{\prime}\in C^{\infty}_{c}(\mathbb{R}^{n})$ be such that $r(D)=r(D)r^{\prime}(D)$. Since $r^{\prime}(D)$ is smoothing, Young’s inequality and (2.2) yield
@@
$\lesssim\int_{\mathbb{R}^{n}}\Big(\int_{\mathbb{R}^{n}}\langle x-y\rangle^{-n-1/2}|(r^{\prime}(D)f)(y)|\mathrm{d}y\Big)^{p}\mathrm{d}x\lesssim\|r^{\prime}(D)f\|_{L^{p}(\mathbb{R}^{n})}^{p}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}^{p}.$
In the remainder, it thus suffices to show that
@@
The high frequencies
Note that
@@
Conclusion of the proof
It remains to extend (3.1) to general $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Here it is relevant to note that, a priori, it is not even clear why the left-hand side of (3.1) is well defined for such $f$. Indeed, initially $e^{it\phi(D)}a(tD)f$ is defined as a tempered distribution for each $t\in I$, but Proposition 2.2 shows that it can be identified with an element of $L^{p}(\mathbb{R}^{n})$. We will prove that there exists a set $U\subseteq\mathbb{R}^{n}$ of measure zero with the property that, for every $t\in I$, there exists a representative $F(\cdot,t)$ of $e^{it\phi(D)}a(tD)f$ in $L^{p}(\mathbb{R}^{n})$ such that the function $x\mapsto\sup_{t\in I}|F(x,t)|\mathbf{1}_{\mathbb{R}^{n}\setminus U}(x)$ is measurable and satisfies $\|\sup_{t\in I}|F(\cdot,t)|\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$.
@@
(3.5) $\big\|\sup_{t\in I}|F(\cdot,t)|\big\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})},$
and
$\big\|\sup_{t\in I}|F_{k}(\cdot,t)-F(\cdot,t)|\big\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f_{k}-f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}.$
In particular, due to the latter inequality, we may also suppose that there exists a set $U\subseteq\mathbb{R}^{n}$ of measure zero such that, for all $x\notin U$, one has $F_{k}(x,t)\to F(x,t)$, uniformly in $t\in I$.
On the other hand, by Proposition 2.2 (1), for each $t\in I$ one has $F_{k}(\cdot,t)\to e^{it\phi(D)}a(tD)f(\cdot)$ in $L^{p}(\mathbb{R}^{n})$ as $k\to\infty$. So there exists a set $U_{t}\subseteq\mathbb{R}^{n}$ of measure zero, and a subsequence $(F_{m_{k,t}})_{k=0}^{\infty}$ of $(F_{k})_{k=0}^{\infty}$, such that $F_{m_{k,t}}(x,t)\to e^{it\phi(D)}a(tD)f(x)$ for $x\notin U_{t}$. It follows that $e^{it\phi(D)}a(tD)f(x)=F(x,t)$ for $x\notin U\cup U_{t}$. Since $e^{it\phi(D)}a(tD)f$ is only determined up to sets of measure zero, $F(\cdot,t)$ is in fact a representative of $e^{it\phi(D)}a(tD)f$ in $L^{p}(\mathbb{R}^{n})$. Now the proof is concluded by applying (3.5). ∎
@@
for any compact interval $I\subseteq(0,\infty)$. Moreover, if $f\in\overline{\mathcal{S}(\mathbb{R}^{n})}\subseteq\mathcal{H}^{s,\infty}_{FIO}(\mathbb{R}^{n})$, then $t\mapsto e^{it\phi(D)}a(tD)f$ is continuous as an $L^{\infty}(\mathbb{R}^{n})$-valued map, and (3.6) holds as well. Finally, by Theorem 4.2, (3.7) is essentially sharp under the same assumptions as mentioned below Theorem 3.2.
The case $p=\infty$ of the other maximal function bounds in this section is straightforward for similar reasons.
@@
The proof is almost identical to that of Theorem 3.2. Firstly, (3.9) follows from (3.8), and it suffices to establish (3.8) for $f\in\mathcal{S}(\mathbb{R}^{n})$, just as in the proof of Theorem 3.2. For such $f$ the left-hand side of (3.8) is well-defined.
Now, to prove (3.8), one first considers the low-frequency component $T_{t}r(D)$ of each $T_{t}$, for $r\in C^{\infty}_{c}(\mathbb{R}^{n})$ satisfying $r(\eta)=1$ if $|\eta|\leq 1/2$, and $r(\eta)=0$ if $|\eta|>1$. Write
$T_{t}r(D)f=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}e^{i(x-y)\cdot\eta}e^{i(\Phi(x,t,\eta)-x\cdot\eta)}a(x,t,\eta)r(\eta)\mathrm{d}\eta f(y)\mathrm{d}y,$
and note that $(x,\eta)\mapsto e^{i(\Phi(x,t,\eta)-x\cdot\eta)}a(x,t,\eta)r(\eta)$ satisfies the conditions of Lemma 3.1, uniformly in $t\in I$. Hence one can apply Young’s inequality and (2.2) to this term, in the same manner as in the proof of Theorem 3.2.
On the other hand, the arguments from the proof of Theorem 3.2 and the bounds in Proposition 2.3 can be combined to deal with the high frequencies. ∎
@@ 3.2. Hypersurfaces
In this subsection we use Theorems 3.2 and 3.5 to prove results about maximal functions associated with hypersurfaces in $\mathbb{R}^{n}$.
For $\Sigma\subseteq\mathbb{R}^{n}$ a (smooth) hypersurface, $\psi\in C^{\infty}_{c}(\mathbb{R}^{n})$, $f\in C(\mathbb{R}^{n})$ and $x\in\mathbb{R}^{n}$, set
@@
Here and throughout, $\mathrm{d}\sigma$ is the induced Lebesgue measure on $\Sigma$, defined as in [40, Section XI.3.1.2] (see also (3.14)). The following result contains Theorem 1.1 as a special case. By Theorem 4.8, it is essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$ whenever $\psi$ is not identically zero.
@@
Let $\Sigma\subseteq\mathbb{R}^{n}$ be a hypersurface with non-vanishing Gaussian curvature, and let $\psi\in C^{\infty}_{c}(\mathbb{R}^{n})$. Then, for all $p\in[1,\infty)$ and $s>d(p)+\frac{1}{p}-\frac{n-1}{2}$, there exists a $C\geq 0$ such that
@@
Fix $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Then the left-hand side of (3.13) is well defined. Moreover, by integration $\mathcal{A}_{t}f$ defines a distribution, i.e. a functional on $C^{\infty}_{c}(\mathbb{R}^{n})$, for every $t\in[1,2]$. We claim that $\mathcal{A}_{t}f=b(tD)f$, where $b:=\widehat{\psi\mathrm{d}\sigma}\in C^{\infty}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ and $b(tD)f$ is initially defined as a distribution as well.
By dilation, it suffices to prove this claim for $t=1$. By definition, one then has
(3.14) $\mathcal{A}_{1}f(x)=\lim_{\varepsilon\to 0}\frac{1}{2\varepsilon}\int_{\mathbb{R}^{n}}f(y)\mathbf{1}_{S_{\varepsilon}}(x-y)\psi(x-y)\mathrm{d}y$
for all $x\in\mathbb{R}^{n}$, where $S_{\varepsilon}:=\{y\in\mathbb{R}^{n}\mid\inf_{z\in\Sigma}|z-y|<\varepsilon\}$. Hence, if we set $\widetilde{g}(x):=g(-x)$ for $g\in C^{\infty}_{c}(\mathbb{R}^{n})$, then
$\int_{\mathbb{R}^{n}}\mathcal{A}_{1}f(x)g(x)\mathrm{d}x$ $=\int_{\mathbb{R}^{n}}\lim_{k\to\infty}\frac{k}{2}\int_{\mathbb{R}^{n}}f(y)\mathbf{1}_{S_{1/k}}(x-y)\psi(x-y)g(x)\mathrm{d}y\mathrm{d}x$
$=\lim_{k\to\infty}\frac{k}{2}\int_{\mathbb{R}^{n}}f(y)\int_{\mathbb{R}^{n}}\mathbf{1}_{S_{1/k}}(x-y)\psi(x-y)g(x)\mathrm{d}x\mathrm{d}y$
$=\int_{\mathbb{R}^{n}}f(y)\lim_{k\to\infty}\frac{k}{2}\int_{\mathbb{R}^{n}}\mathbf{1}_{S_{1/k}}(x-y)\psi(x-y)g(x)\mathrm{d}x\mathrm{d}y$
$=\int_{\mathbb{R}^{n}}f(-y)b(D)\widetilde{g}(y)\mathrm{d}y,$
and the latter is the adjoint action of the distribution $b(D)f$ on $g$. Note that, in this chain of identities, we used both the dominated convergence theorem and Fubini’s theorem. This is in turn allowed because $f$ is locally bounded and, given that $\psi$ and $g$ are compactly supported, the variables xyx-y and $x$ are restricted to compact sets, meaning that $y$ is as well. This proves the claim.
Now, we will show that $b(tD)f\in L^{p}(\mathbb{R}^{n})$ for every $t\in[1,2]$, and that
(3.15) $\big\|\sup_{1\leq t\leq 2}|b(tD)f|\big\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}.$
It then follows that $\mathcal{A}_{t}f\in L^{p}(\mathbb{R}^{n})$ as well, and that (3.13) holds, as required.
After decomposing the support of $\psi$ into finitely many pieces and changing coordinates, we may suppose that
(3.16) $b(\xi)=a(\xi)e^{i\phi(\xi)}+e(\xi)$
for all $|\xi|\geq 1$ (see e.g. [40, Section VIII.5.7] or [37, Theorem 1.2.1]). Here $\phi\in C^{\infty}(V)$ is real-valued and positively homogeneous of degree $1$, defined on an open conic subset $V\subseteq\mathbb{R}^{n}\setminus\{0\}$, and $\mathrm{rank}\,\partial_{\xi\xi}^{2}\phi(\xi)=n-1$ for all $\xi\in V$. Also, $a\in S^{-(n-1)/2}(\mathbb{R}^{n})$ satisfies $\mathrm{supp}(a)\subseteq V^{\prime}\cap\{\xi\in\mathbb{R}^{n}\mid|\xi|\geq 1/2\}$ for some conic $V^{\prime}\subseteq V$ which is closed in $\mathbb{R}^{n}\setminus\{0\}$, and $e\in S^{m}(\mathbb{R}^{n})$ for all $m\in\mathbb{R}$. In fact, given that the low-frequency component of $b$ is itself an element of $S^{m}(\mathbb{R}^{n})$ for all $m\in\mathbb{R}$, (3.16) holds for all $\xi\in\mathbb{R}^{n}$.
By Proposition 2.2 and (3.16), $e^{it\phi(D)}a(tD)f$, $e(tD)f$ and $b(tD)f$ are all elements of $L^{p}(\mathbb{R}^{n})$, for every $t\in[1,2]$. Moreover,
$\sup_{1\leq t\leq 2}|b(tD)f(x)|\leq\sup_{1\leq t\leq 2}\left(|e^{it\phi(D)}a(tD)f(x)|+|e(tD)f(x)|\right)$
for almost all $x\in\mathbb{R}^{n}$. Next, for each $s>d(p)+\frac{1}{p}-\frac{n-1}{2}$, Theorem 3.2 yields
@@
On the other hand, by (3.10) (with $\phi\equiv 0$ and $m=d(p)-s(p)-\frac{n-1}{2}$), one has